### v1

# Evidence of partonic collectivity in ultra-relativistic heavy-ion collisions with NCQ scaling of radial flow

Abstract
We report the first observation of Number of Constituent Quark (NCQ) scaling of the radial flow observable $v_{0}(p_{\rm T})$ in relativistic heavy-ion collisions. Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV are investigated using the string-melting version of the AMPT model, while Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV are analyzed using published ALICE data, with calculations from PYTHIA8/Angantyr serving as a non-collective baseline. The $v_{0}(p_{\rm T})$ spectra from AMPT-SM exhibit clear collective signatures, including mass ordering at lower-$p_{\rm T}$ and a meson–baryon separation at intermediate-$p_{\rm T}$, which are absent in the baseline calculations. In central collisions, $v_{0}(p_{\rm T})/n_{q}$ follows robust NCQ scaling when expressed as a function of transverse kinetic energy per quark, $(m_{\rm T}-m_{0})/n_{q}$, while significant deviations emerge toward peripheral events, indicating a stronger manifestation of collectivity in central collisions. The scaling is more precise at RHIC than at LHC energies, consistent with earlier observations for elliptic flow $v_{2}$. These findings provide strong evidence that radial collectivity is established predominantly at the partonic stage of the fireball evolution.
keywords:
QGP , radial flow , NCQ scaling , transport model , PYTHIA8/Angantyr
journal: Physics Letters B
[…]
[label1]organization=Department of Physics, Bodoland University, addressline=Rangalikhata, city=Kokrajhar, postcode=783370, state=Assam, country=India
[…]
[label2]organization=Department of Physics, Kokrajhar University, city=Kokrajhar, postcode=783370, state=Assam, country=India
1 Introduction
Ultra-relativistic nucleus-nucleus collisions provide access to hot and dense QCD matter, commonly known as the quark-gluon plasma (QGP), where quarks are no longer confined within hadrons [1, 2, 3, 4, 5]. The formation of this strongly interacting plasma is now supported by a broad set of experimental observations. Two of its most compelling signatures are jet quenching [6, 7] and the Number of Constituent Quark (NCQ) scaling of the second-order flow harmonic $v_{2}$ [8, 9]. The observed scaling of elliptic flow of hadrons, collectively follow a universal curve when scaled by their number of constituent quarks, provides direct evidence for the emergence of partonic collectivity prior to hadronization.
Complementary to elliptic flow, the isotropic radial-flow originates from the strong pressure gradients generated at the early stage of the collision, driving the collective transverse expansion of the medium and providing insight into the equation of state and space-time evolution of the QGP [10]. Traditionally, radial flow has been characterized using a transverse momentum ($p_{\rm T}$)-integrated observable, obtained by performing simultaneous blast-wave fits to the $p_{\rm T}$-spectra of identified hadrons. Recently, a new observable, $v_{0}(p_{\mathrm{T}})$, has been introduced to quantify radial flow through the correlation between the particle yield in a given $p_{\mathrm{T}}$-bin and the eventwise mean transverse momentum fluctuation [11, 12]. This differential measure provides a more direct handle on collective expansion and exhibits improved sensitivity compared to traditional blast-wave-based extractions that rely on model-dependent assumptions [13, 14]. Recent measurements by ALICE collaboration [10] on $v_{0}(p_{\mathrm{T}})$ exhibit a clear mass ordering at low $p_{\mathrm{T}}$ and a meson-baryon splitting at intermediate $p_{\mathrm{T}}$, consistent with the qualitative behavior previously observed in $v_{2}$ analyses [15, 16]. The observable is also sensitive to the transport properties of the system, with a pronounced dependence on the bulk viscosity [10, 17], in contrast to elliptic flow, whose magnitude is influenced by both bulk and shear viscous contributions [18, 19]. Inclusive charged-particle measurements demonstrate that the radial-flow observable $v_{0}(p_{\mathrm{T}})$ exhibits genuine collective behavior, manifested through long-range pseudorapidity correlations, factorization of two-particle correlations into single-particle $v_{0}(p_{\mathrm{T}})$, and a nearly centrality-independent normalized shape $v_{0}(p_{\mathrm{T}})/v_{0}$ at low $p_{\mathrm{T}}$ [20, 21]. Further, as reported in Refs. [22, 14], the hydrodynamic-inspired blast-wave model with fluctuating kinetic freeze-out temperature and flow velocity can explain the mass-ordering of $v_{0}(p_{\rm T})$ for identified charged hadrons at low-$p_{\rm T}$. Moreover, as an extension, an AMPT-based study [22] with string melting configuration demonstrated that the meson-baryon crossing at intermediate-$p_{\rm T}$ originates primarily from the quark coalescence mechanism.
A recent investigation of elliptic flow at lower collision energies has established that signatures such as mass ordering can arise even in systems that do not reach full local thermalization. In particular, the study at $\sqrt{s_{NN}}=4.5$ GeV reported in Ref. [9] demonstrated that mass ordering may originate, at least in part, from hadronic interactions during the late stages of the evolution. Such findings introduce an important ambiguity regarding the interpretation of the LHC measurements: the observed collectivity in $v_{0}(p_{\mathrm{T}})$ may receive contributions from both the partonic and hadronic phases.
This motivates a systematic investigation of the dynamical origin of the $v_{0}(p_{\mathrm{T}})$ signal. In the present work, we examine its scaling properties with the number of constituent quarks for the first time within the AMPT String Melting (SM) framework, which provides controlled access to partonic and hadronic transport effects. The results are compared with existing experimental measurements to assess whether the observed behavior can be attributed predominantly to partonic collectivity or can also emerge from hadronic dynamics.
2 Model Description
2.1 The AMPT model
The multi-phase transport (AMPT) model provides a comprehensive microscopic description of heavy-ion collisions and consists of four distinct stages [23]. In the initial stage, HIJING [24, 25] generates the fluctuating nucleon-nucleon interactions, including hard and semi-hard scatterings, jet production, nuclear shadowing, and the associated excited strings. In the default version, these strings hadronize via Lund string fragmentation, whereas in the string-melting (SM) version, they are first converted to their constituent quarks and antiquarks according to the endpoint flavors and momenta, providing a more suitable setup for studying partonic collectivity. The subsequent partonic evolution (present only in the string-melting scenario) is modeled with Zhang’s Parton Cascade (ZPC[26], which includes elastic two-body scatterings among quarks and gluons with parameterized cross-sections, followed by hadronization through a quark coalescence prescription in which nearby partons recombine into mesons and baryons. Finally, the hadronic phase is described by the ART (A Relativistic Transport) model [27], which propagates the produced hadrons through elastic and inelastic meson–meson, meson–baryon, and baryon-baryon interactions, including resonance formation and decay channels. For the present study, we employ the AMPT model version 2.26t9b.
2.2 The PYTHIA8/Angantyr model
The Angantyr model [28] is an extension of the PYTHIA8 framework [29] specifically designed to simulate ppA and AA collisions by stacking multiple fluctuating nucleon-nucleon sub-collisions on top of the standard $pp$ framework, without introducing an explicit hydrodynamic medium or QGP. In Angantyr, the nuclear geometry is determined through a Glauber model calculation that incorporates event-by-event fluctuations of the nucleon–nucleon cross-section. Each binary encounter is subsequently classified as non-diffractive, single-diffractive, or double-diffractive excitation, following a wounded-nucleon-based approach inspired by the Fritiof model [30]. Multi-parton interactions and color reconnection are handled by the underlying PYTHIA8 machinery, which allows Angantyr to generate fully exclusive hadronic final states and to describe global observables such as charged-particle multiplicity and transverse-momentum spectra in heavy-ion collisions [31, 32, 33]. The Angantyr model serves as a non-collective baseline for studies such as radial flow, by quantifying the contributions from initial-state fluctuations and final-state partonic interactions that are not associated with a thermalized medium. The Pb+Pb collision events at $\sqrt{s_{\rm NN}}$ = 5.02 TeV for this analysis are generated using this model.
3 Methodology
3.1 Event Selection and Centrality Determination
In the present analysis, heavy-ion collisions have been simulated for two collision systems. For Au+Au collisions at RHIC energy ($\sqrt{s_{\rm NN}}$ = 200 GeV), a total of 19.9 million minimum-bias events were produced using the AMPT model in its String Melting configuration. For Pb+Pb collisions at LHC energy ($\sqrt{s_{\rm NN}}$ = 5.02 TeV), 3.4 million minimum-bias events were generated using the PYTHIA8/Angantyr model. Centrality, which serves as a proxy for the collision geometry and overlap region, is determined from final-state charged-particle multiplicity. The method is chosen to align with established experimental procedures for each energy regime. In the Au+Au system, centrality for the AMPT-SM events is determined by the charged-particle multiplicity at mid-rapidity ($|\eta|\leq 0.5$), consistent with the standard procedure employed in analyses of RHIC data. For the Pb+Pb system, centrality in the PYTHIA8/Angantyr model is classified using the total charged-particle multiplicity within the pseudorapidity acceptance of the ALICE V0 detector ($2.8<\eta<5.1$ and $3.7<\eta<-1.7$), a method denoted as V0M. This directly mirrors the experimental centrality determination used by ALICE in the measurement of $v_{0}(p_{\rm T})$ [10]. From these samples, specific centrality percentiles are selected for detailed study: 0–20% (most central) and 20–40% (semi-central) for Au+Au collisions, and 10–20%, 30–40%, and 60–70% for Pb+Pb collisions.
Figure 1: (Color online) Transverse momentum ($p_{\rm T}$) dependence of the radial flow observable $v_{0}(p_{\rm T})$ for identified hadrons in Au+Au collisions generated with the String Melting version of AMPT model, shown for 0-20% (most central) and 20-40% (semi-central) centrality selections. Shaded bands indicate bootstrap-estimated statistical uncertainties.
3.2 Observable
For the present investigation, radial flow is estimated using the novel observable, $v_{0}(p_{\rm T})$, which provides a unique probe of long-range collective dynamics by quantifying event-by-event correlations between fluctuations in charged-particle multiplicity and fluctuations in the average transverse momentum [11]. The observable $v_{0}(p_{\rm T})$ is defined as,
$v_{0}(p_{\rm T})=\frac{\langle F_{\rm A}^{p_{\rm T}}{[p_{\rm T}]}_{\rm B}\rangle-\langle F_{\rm A}^{p_{\rm T}}\rangle{\langle[p_{\rm T}]}_{\rm B}\rangle}{\langle F_{\rm A}^{p_{\rm T}}\rangle\sigma_{[p_{\rm T}]}},$ (1)
where, $F_{\rm A}^{p_{\rm T}}$ is the fraction of particles in the transverse momentum bin $p_{\rm T}$ within the pseudorapidity window A and is given by,
$F_{\rm A}^{p_{\rm T}}=\frac{N_{\rm ch}^{\rm A}(p_{\rm T})}{N_{\rm ch}^{\rm A}},$ (2)
where $N_{\rm ch}^{\rm A}(p_{\rm T})$ is the number of particles in the $p_{\rm T}$ bin and $N_{\rm ch}^{\rm A}$ is the total multiplicity in window A.The second, correlating quantity is the event-averaged mean transverse momentum, measured in a separate, longitudinally-displaced window B (illustrated in Fig. 2), with a gap, $\eta_{\rm gap}>0$ to suppress short-range correlations, is given by,
${[p_{\rm T}]}_{\rm B}=\frac{1}{N_{\rm ch}^{\rm B}}\sum_{i\in B}p_{\rm T_{i}}$ (3)
{tikzpicture}
[…]
Figure 2: Cartoon illustrating the $\eta$ regions used in the analysis. Regions A (red) and B (green) are separated by a gap $\eta_{\rm gap}$. The observable $F_{\rm A}^{p_{\rm T}}$ is correlated with the transverse momentum $p_{\rm T}]_{\rm B}$ measured in the opposite region.
The term $\sigma_{[p_{\rm T}]}$ in Eqn. 1 can be expressed as follows,
$\sigma_{[p_{\rm T}]}=\sqrt{\langle{[p_{\rm T}]}_{\rm A}{[p_{\rm T}]}_{\rm B}\rangle-\langle{[p_{\rm T}]}_{\rm A}\rangle\langle{[p_{\rm T}]}_{\rm B}\rangle}.$ (4)
Here, $\langle..\rangle$ represents event-average and $\sigma_{[p_{\rm T}]}$ is the standard deviation of mean-$p_{\rm T}$ fluctuations.
Figure 3: (Color online) The $v_{0}(p_{\rm T})$ observable as a function of transverse kinetic energy $\rm(m_{T}-m_{0})$ for identified hadrons in AMPT-SM simulations of Au+Au collisions, displayed for 0-20% (most central) and 20-40% (semi-central) centrality bins. Bootstrap-derived statistical uncertainties are shown as shaded regions.
Figure 4: (Color online) Comparison of $v_{0}(p_{\rm T})$ between PYTHIA8/Angantyr predictions and ALICE measurements [10] for identified hadrons in Pb+Pb collisions at = $\sqrt{s_{\rm NN}}$ = 5.02 TeV, presented for 10-20% (left panel) and 30-40% (right panel) centrality classes. Model calculations are represented by solid curves with shaded bands denoting bootstrap uncertainties, while experimental data are shown as filled markers.
Figure 5: (Color online) $v_{0}(p_{\rm T})$ of identified hadrons presented in terms of transverse kinetic energy, ($m_{\rm T}-m_{\rm 0}$) for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. The plot is compiled from the experimental data of ALICE [10].
Figure 6: (Color online) Number of Constituent Quark (NCQ) scaled radial flow observable, $v_{0}(p_{\rm T})/n_{q}$, versus $p_{\rm T}/n_{q}$ (top panel) and $(m_{T}-m_{0})/n_{q}$ (bottom panel) for identified hadrons in AMPT-SM simulations of Au+Au collisions at 0-20% (most central) and 20-40% (semi-central) centrality. A third-order polynomial fit to $\pi^{\pm}$ data (solid curve) serves as the NCQ scaling reference. Bootstrap uncertainties are displayed as vertical error bars.
Figure 7: (Color online) NCQ-scaled $v_{0}(p_{\rm T})$ of identified hadrons as a function of NCQ-scaled $p_{\rm T}$ (top panel) and as a function of NCQ-scaled ($m_{\rm T}-m_{\rm 0}$) (bottom panel) in Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. This plot is compiled using the experimental data of ALICE [10].
4 Results and Discussion
4.1 Estimation of $v_{0}(p_{\rm T})$ at $\sqrt{s_{\rm NN}}=200$ $\rm GeV$
In Fig. 1, we display the radial flow observable, $v_{0}(p_{\rm T})$ for identified hadrons ($\pi^{\pm}$, $K^{\pm}$, and $p+\overline{p}$) in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV, obtained from AMPT String Melting simulations for central (0-20%) and semi-central (20-40%) class. The observable is calculated using the subevent method with $\eta_{\rm gap}$ = 0.4 to minimize the short-range non-flow correlations. A clear mass ordering is observed at lower-$p_{\rm T}$ ($p_{\rm T}\leq 2$ GeV/c), while a characteristic baryon–meson crossover emerges at intermediate $p_{\rm T}$ ($2\leq p_{\rm T}\leq 3$ GeV/c). These features are consistent with expectations from collective radial expansion of the produced medium. The trends align with measurements at $\sqrt{s_{\rm NN}}=5.02$ TeV by ALICE [10] and with hydrodynamic model calculations [17].
[…]
4.2 Estimation of $v_{0}(p_{\rm T})$ at $\sqrt{s_{\rm NN}}=5.02$ TeV
This analysis is extended to Pb+Pb collisions at $\sqrt{s_{\rm NN}}$ = 5.02 TeV by comparing the radial flow observable $v_{0}(p_{\rm T})$ for identified charged hadrons from PYTHIA8/Angantyr simulations with the corresponding ALICE data [10] for 10–20% (central) and 30–40% (semi-central) classes, as shown in Fig. 4. As reported by the ALICE collaboration [10], the experimental results display clear signatures of collective expansion, characterized by a pronounced mass ordering at low transverse momentum ($p_{\rm T}\leq 3~\text{GeV}/\textit{c}$) and a distinct meson-baryon crossing at intermediate transverse momentum ($p_{\rm T}\geq 3~\text{GeV}/\textit{c}$). In contrast, the $v_{0}(p_{\rm T})$ values predicted by PYTHIA8/Angantyr remain close to zero over the entire $p_{\rm T}$ range, exhibiting neither mass ordering nor meson-baryon crossing. This pronounced discrepancy reflects the absence of a hydrodynamic evolution stage in the model, which prevents the buildup of sustained pressure gradients required to generate collective radial flow. The inability of the model to reproduce the observed experimental features provides further evidence that the measured $v_{0}(p_{\rm T})$ patterns arise from genuine final-state collective dynamics, rather than from initial-state effects or string-fragmentation mechanisms.
[…]
4.3 NCQ scaling of radial flow observable $v_{0}(p_{\rm T})$
4.3.1 NCQ scaling at $\sqrt{s_{\rm NN}}=200$ $\rm GeV$
We first apply Number of Constituent Quark (NCQ) scaling to the $v_{0}(p_{\rm T})$ observable by rescaling both the flow magnitude and the kinematic variables with the number of constituent quarks $n_{q}$. Fig. 6 presents the resulting scaled distributions, $v_{0}(p_{\rm T})/n_{q}$, for Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV simulated using the AMPT-SM model. The results are shown as functions of $p_{\rm T}/n_{q}$ (top) and $(m_{\rm T}-m_{0})/n_{q}$ (bottom) for central (0–20%) and semi-central (20–40%) events. While the scaling with $p_{\rm T}/n_{q}$ is only approximate, a visibly improved scaling is observed when the transverse kinetic energy variable $(m_{\rm T}-m_{\rm 0})/n_{q}$ is employed. Moreover, the degree of NCQ scaling is found to be systematically better for central (0–20%) collisions compared to the semi-central (20–40%) class, indicating a stronger manifestation of collectivity in more central events.
To quantify the observed scaling behavior, we have adopted the following procedure. The $v_{0}(p_{T})$ spectrum of pions ($\pi^{\pm}$) is fitted with a third-order polynomial (solid line), which is then used as a reference to extract a scaling parameter $S_{f}$. This parameter measures the deviation of the kaon ($K^{\pm}$) and proton ($p+\bar{p}$) $v_{0}(p_{T})$ distributions from the pion reference. $S_{f}$ is therefore defined as,
$S_{f}=\sum_{i=1}^{2}\sum_{j=1}^{N}|y_{j}-f(x_{j})|_{i}$ (5)
where $N$ is the number of bins in which $v_{0}(p_{\rm T})$ is calculated for $K^{\pm}$ and $p+\overline{p}$ and $y_{j}$ correspond to $v_{0}(p_{\rm T})$ values of $K^{\pm}$ and $p+\overline{p}$. By definition, for ideal scaling, $S_{f}$ must be equal to zero.
The extracted $S_{f}$ values for NCQ scaling with respect to $p_{\rm T}/n_{q}$ and $(m_{\rm T}-m_{0})/n_{q}$ for different centrality classes are summarized in Tab. 1. While the qualitative scaling features have already been discussed above, the $S_{f}$ values provide a quantitative validation of these observations. Two clear trends emerge: (i) the scaling improves significantly when $(m_{\rm T}-m_{0})/n_{q}$ is used as the scaling variable, and (ii) the scaling is systematically better for more central collisions.
Table 1: Scaling parameter, $S_{f}$ values for scaled AMPT Au+Au results (top) and scaled ALICE [10] Pb+Pb results (bottom).
$S_{f}$ Values
System $(\sqrt{s_{\rm NN}})$ Centrality $p_{\rm T}/n_{q}$ $(m_{T}-m_{0})/n_{q}$
Au+Au (200 GeV) 0-20% $0.072$ $0.046$
20-40% $0.085$ $0.066$
10-20% $0.300$ $0.132$
Pb+Pb (5.02 TeV) 30-40% $0.373$ $0.186$
60-70% $0.629$ $0.373$
4.3.2 NCQ scaling at $\sqrt{s_{\rm NN}}=5.02$ $\rm TeV$
We extend the NCQ scaling analysis to LHC energies by compiling the published ALICE $v_{0}(p_{\rm T})$ data for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [10]. Fig. 7 presents the scaled distributions $v_{0}(p_{\rm T})/n_{q}$ as functions of $p_{\rm T}/n_{q}$ (top) and $(m_{\rm T}-m_{0})/n_{q}$ (bottom) for different centrality classes. A clear NCQ scaling is observed for central (10–20%) and semi-central (30–40%) collisions when the transverse kinetic energy per quark, $(m_{\rm T}-m_{0})/n_{q}$, is used as the scaling variable. In contrast, the scaling with $p_{\rm T}/n_{q}$ remains visibly weaker for the same centrality classes. For peripheral (60–70%) collisions, the scaling is found to break down for both variables, indicating a reduced degree of partonic collectivity. The violation of NCQ scaling in peripheral events can be ascribed to the smaller system size and shorter lifetime of the produced medium, which limit the buildup of strong collective dynamics and possibly enhance the relative contribution from non-flow effects and initial-state fluctuations. Conversely, the improved universality of the $(m_{\rm T}-m_{0})/n_{q}$–scaled distributions across hadron species in central collisions supports a scenario in which the observed collectivity is established at the partonic level.
To quantify the degree of NCQ scaling at LHC energies, we apply the same fitting and evaluation procedure defined in Sec. 4.3.1, using the scaling parameter $S_{f}$ [Eqn. (5)]. The extracted $S_{f}$ values, summarized in Tab. 1, show that the scaling is systematically stronger for $(m_{\rm T}-m_{0})/n_{q}$ than for $p_{\rm T}/n_{q}$, and improves toward more central collisions, while being substantially degraded in peripheral events.
5 Summary
In summary, this Letter presents the first systematic study of Number of Constituent Quark (NCQ) scaling of the novel radial-flow observable $v_{0}(p_{\rm T})$. In the absence of published RHIC data, Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV are analyzed using the string-melting configuration of the multi-phase transport (AMPT) model, while Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV are investigated using published ALICE data [10], with PYTHIA8/Angantyr employed as a non-collective baseline.
Our main observations can be summarized as follows: (i) The $v_{0}(p_{\rm T})$ spectra obtained from AMPT-SM exhibit characteristic signatures of collective behavior, including mass ordering at lower-$p_{\rm T}$ and a baryon–meson separation at intermediate-$p_{\rm T}$, whereas such features are absent in the PYTHIA8/Angantyr calculations. (ii) A direct comparison with ALICE Pb+Pb data reveals that PYTHIA8/Angantyr fails to reproduce the observed $v_{0}(p_{\rm T})$ systematics, providing strong evidence that the experimentally observed correlations cannot be explained without invoking collective dynamics. (iii) Since mass ordering alone is insufficient to establish the microscopic origin of collectivity, we perform a detailed NCQ-scaling analysis. For AMPT-SM, the $v_{0}(p_{\rm T})$ values for pions, kaons, and protons exhibit robust NCQ scaling in central collisions, while clear violations are observed toward peripheral events. (iv) The scaling is found to be significantly improved when expressed in terms of transverse kinetic energy per quark, $(m_{\rm T}-m_{0})/n_{q}$, compared to $p_{\rm T}/n_{q}$, consistent with observations previously reported for elliptic flow $v_{2}$ [36]. (v) A comparative study further indicates that the NCQ scaling of $v_{0}(p_{\rm T})$ is more precise at RHIC energies than at the LHC, following a hierarchy similar to that established for anisotropic flow.
The observation of NCQ scaling in $v_{0}(p_{\rm T})$ provides compelling evidence that the long-range momentum correlations responsible for radial expansion are established at the partonic level of the fireball evolution. This result extends the paradigm of quark-level collectivity from anisotropic flow observables to isotropic radial dynamics, offering a unified and more complete picture of the nearly perfect fluid behavior of the strongly interacting quark-gluon plasma.